OpenStudy (anonymous):

Find the solution to the system of equations. 3x + y + z = 6 3x - y + 2z = 9 y + z = 3 A. x = 2, y = 1, z = 4 B. x = 2, y = -1, z = 5 C. x = 5, y = -2, z = 6 D. x = 1, y = 0, z = 3

Parth (parthkohli):

x = 1 3 - y + 2z= 9 -y + 2z = 6 D is the answer